OpenStudy (anonymous):

What transformations ate required to obtain f(x) = 3/x+10 from the graph of y =1/x

OpenStudy (jdoe0001):

\(\bf f(x) = \cfrac{3}{x}+10\qquad \qquad g(x)= \cfrac{1}{x}\\ \quad \\ \color{red}{3}\left(\cfrac{1}{x}\right)\color{red}{+10}\implies f(x)\) y = A ( Bx + C ) + D A = vertical stretch/shrink of the graph D = vertical shift, D > 0, up, D < 0, down